OpenStudy (anonymous):

A bag contains blue and yellow marbles. Two marbles are drawn without replacement. The probability of selecting a blue marble and then a yellow marble is 0.37, and the probability of selecting a blue marble on the first draw is 0.55. What is the probability of selecting a yellow marble on the second draw, if the first marble drawn was blue? A) 18% B) 20% C) 46% D) 67%

OpenStudy (anonymous):

D) 67%

OpenStudy (anonymous):

P(B)*P(Y|B)=.37 P(B)=.55

OpenStudy (anonymous):

solve for P(Y|B)

OpenStudy (gw2011):

Let B=Blue Let Y=Yellow (B)(Y)=0.37 B=0.55 (0.55)(Y)=0.37 Y=0.37/0.55 Y=0.67=67% Therefore, the answer is D which is 67%